A car salesperson has a 20% chance of landing a sale with a random customer on the lot. Suppose 12 people come onto the lot today.

What is the standard deviation of cars he or she is going to sell today? (Keep the answer with 2 decimal places) (no excel work)
Answer #1

Solution

Given that,

p = 0.20

1 - p = 1 -0.20 =0.80

n = 12

= p =0.20

standard deviation =  [p( 1 - p ) / n] = [(0.20*0.80) /12 ] = 0.12